10-x+3x=2x-7+5x
solving for the unknown
not sure of the steps to solve this one because the varible is on both sides
I was given this question and asked to write it down from the board

Answer by stanbon(75887)

10-x+3x=2x-7+5x
Simplify each side:
10+2x = 7x-7
Add 7 to both sides to get:
17+2x = 7x
Subtract 2x from both sides:
17 = 5x
Divide both sides by 5 to get:
x = 17/5
-----------------
Cheers,
Stan H.